OpenStudy (anonymous):

Suppose Q=x^1y^4. If dQ/dt=-2 and dx/dt=3, find dy/dt when x=1 and y=1

OpenStudy (anonymous):

Implicit differentiation: \[Q=xy^4~~\Rightarrow~~\frac{dQ}{dt}=y^4\frac{dx}{dt}+4xy^3\frac{dy}{dt}\] Plug in the given info.

OpenStudy (anonymous):

i am getting -2/7. It is coming out wrong?

OpenStudy (anonymous):

I am putting -2=(1)^4 times 3+4(1)(1)^3dy/dt

OpenStudy (anonymous):

Maybe I am putting it in the calculator wrong

OpenStudy (anonymous):

\[-2=(1)^4(3)+4(1)(1)^3\frac{dy}{dt}~~\Rightarrow~~-2=3+4\frac{dy}{dt}~~\Rightarrow~~\frac{dy}{dt}=-\frac{5}{4}\]

OpenStudy (anonymous):

Got it. Thanks

OpenStudy (anonymous):

yw